OpenStudy (anonymous):

A person's blood pressure P at time t (in seconds) is given by P = 100 + 20cost(6t). Find the maximum value of P (called the systolic pressure) and the minimum value of P (called the diastolic pressure) and give one or two values of t where these maximum and minimum values of P occur.

OpenStudy (cwrw238):

first find the derivative dP/dt and equate it to zero, then solve for t

OpenStudy (anonymous):

So I found the derivative: -120sin6t

OpenStudy (anonymous):

So when I set it equal to 0: t = 0, pi/6, pi/3, pi/2... Now what

OpenStudy (cwrw238):

thats correct so -120sin(6t) = 0

OpenStudy (cwrw238):

now you need to find which values of t give maximum and minimum values of P you can do this by using the second derivative test

OpenStudy (anonymous):

So I've derived it a second time and got -720cos6t

OpenStudy (anonymous):

I'm just not really sure how to go further from here.

OpenStudy (cwrw238):

ok now plug in , say, t = pi/6 and see what you get

OpenStudy (anonymous):

I think I get 720

OpenStudy (cwrw238):

yes as cos pi = -1 a positive 2nd derivative indicates a minimum

OpenStudy (anonymous):

so at pi/6 the function is a minimum?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

and how did you just randomly choose pi/6?

OpenStudy (cwrw238):

t = pi/3 gives -720 cos 6pi/ 3

OpenStudy (cwrw238):

why not? the question asked for one or two values of t givibf min or max

OpenStudy (cwrw238):

lol * giving

OpenStudy (cwrw238):

pi/3 will give a maximum value as second derivative = -720

OpenStudy (anonymous):

So do I just keep doing that for all the values?

OpenStudy (anonymous):

0, pi/6, pi/3, pi/2? and so on?

OpenStudy (cwrw238):

the original function is periodic so there are many peaks and troughs

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

at what value do I stop though?

OpenStudy (cwrw238):

you need only find 2 values for P :- plug in pi/6 and pi/3 into the original function

OpenStudy (anonymous):

wait, why don't I find values for 0, pi/2?

OpenStudy (cwrw238):

eg minimum p = 100 + 20 cos 6* PI/6

OpenStudy (cwrw238):

you can find values for 0 and pi/2 if you want to but the question asks for 1 Or 2

OpenStudy (anonymous):

okay i see Thank you!

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

you should get diastolic 80 and systolic = 120

OpenStudy (anonymous):

Got it!